OpenStudy (inspectorjoe):

Need help please 27x^6-8 factoring the sum or difference of cubes

OpenStudy (inspectorjoe):

\[27x ^{6}-8\]

OpenStudy (inspectorjoe):

keep in mind I'm factoring the sum or deference of cubues

OpenStudy (madhu.mukherjee.946):

what is your question

OpenStudy (inspectorjoe):

27x^6-8 factoring the sum or difference of cubes

OpenStudy (madhu.mukherjee.946):

(3x^2)^3-2^3

OpenStudy (inspectorjoe):

how did we get there?

OpenStudy (zzr0ck3r):

\[27x^6-8=3^3(x^2)^3-2^3=(3x^2)^3-2^3\] Now we use the difference of cubes formula that states \(a^3-b^3=(a-b)(a^2+ab+b^2)\)

OpenStudy (zzr0ck3r):

Can you please delete the huge space post?

OpenStudy (inspectorjoe):

howhow?

OpenStudy (inspectorjoe):

got it thanks

OpenStudy (inspectorjoe):

and thanks for the help